LHC commissioning - latest news

Week 51 - Machine coordinators: Roger Bailey and Gianluigi Arduini

Last update:Thu, 17-Dec-2009 13:49

Day

Short summary

Link to report from 8:30

Monday

14th December

Problem on current leads Q10.L4.B2 needing access

Also exchange vacuum pump for S34 leak pumping

Then access needed for beam dump power supply

Ready to inject at 18.00

Setting up 16 on 16

RT trims from tune feedback

  ALICE solenoid OFF

    Separation bumps ON

    Looked good for lifetime initially but &

    Inject 16 on 16 into separation bumps (off in IR 8 backgr)

    Bumps off one by one

    Experiments check collision points OK

16 on 16 fill

    ALICE solenoid OFF for 1 hour

8:30 Summary slides for December 14

Tuesday

15th December

Beam lost around 02:00 - vacuum pressure rise understood - more details here

  • 08:30-10:30 TOTEM
  • 13:00-15:00 Abort gap cleaning
  • 14:00-19:00 Orbit feedback
  • 19:00 preparation for ramp & squeeze
  • 21:30 Successful ramp. Beam dumped due to trip of triplet orbit corrector as a result of a "wild" trim from real-time channel
  • 21:30-01:00 Pre-cycle

 

8:30 Summary slides for December 15

Wednesday

16th December

  • 01:00-02:30 Second successful ramp. Optics measurements
  • 02:30-03:30 Squeeze in IR5 to 9 and 7 m - Optics measurements
  • 03:30-04:30 Hands-off time
  • 04:30-05:30 Optics measurements with squeezed optics
  • 05:30-06:00 Energy extraction test on bends of Sector 12
  • 06:00-08:00 Pre-cycle
  • 08:30 Current lead control problem on RQTL11.L4B1
  • 09:38 RQ9.R8B2 quench heater fired (likely instrumentation problem)
  • 10:30-12:00 Precycling
  • 12:00-14:00 ATLAS aperture
  • 14:00-15:00 Injection studies
  • 15:00-16:00 Lifetime studies
  • 16:00-16:30 studies on 8 kHz noise
  • 16:30-17:00 BLM-nQPS test
  • 17:00-18:00 Intrusion test
  • 18:00 End of run

8:30 Summary slides for December 16 (pptx)